From the top of a cliff, Tony Stark can see a boat on the ocean. The angle of depression is 41°. If Tony Stark uses his laser range finder and determines that the boat is 1.4 kilometers away, how high is the cliff?


1
Expert's answer
2021-02-24T06:09:29-0500

"\sin 41^o = \dfrac{opposite}{hypotenuse}\\
\sin 41^o = \dfrac{h}{1.4km}\\
h = 1.4km \times sin 41^o\\
h = 1.4km \times 0.6561\\
h = 0.91854 km\\
\implies \bold {h = 918.54 metres}"